# Impact of evanescent scattering modes and finite dispersion in superconducting junctions

### v1

@@ III.1 Detaching from continuum
To set the stage, let us first consider a generic case of an energy-dependent scattering matrix, and show how this energy-dependence is noticeable within the ABS spectrum, at least within the above described standard literature formalism.
For this purpose, we focus in particular on the previously discussed fact that the entire ABS energy spectrum approaches the gap $\Delta$ for zero phase differences, ${\widehat{\phi}=0}$, and thus touches the continuum. As already explained, for a symmetric scattering matrix, the origin of the ${E=\Delta}$ solution can be conveniently seen on the level of Eq. (18), where for ${E=\Delta}$ and ${\widehat{\phi}=0}$, ${\alpha^{2}S_{e}e^{i\widehat{\phi}}S_{h}e^{-i\widehat{\phi}}=S_{e}S_{e}^{*}=1}$. If we now include a finite energy dependence in the scattering, we instead get the term $S_{e}(E)S_{e}^{*}(-E)$ (due to particle-hole symmetry), which in general no longer reduces to $1$. As a consequence, the ABS spectrum detaches from the continuum at $\widehat{\phi}=0$.
@@
The second peculiarity is that it does not return the eigenenergies directly, but instead, it provides what we refer to as the pseudoenergy $\widetilde{E}=\sqrt{\Delta^{2}-E^{2}}$. This is because the subtraction of Eq. (23) from Eq. (22) yields $\text{Im}[\alpha]=-\sqrt{1-E^{2}/\Delta^{2}}$ on the right hand side. Obviously, the addition of Eqs. (22) and (23) (which was the original procedure of Ref. [13]) would instead yield $\text{Re}[\alpha]=E/\Delta$ and would thus be much closer to the generally more conventional form of an effective Schrödinger equation. But, the above chosen form has a central advantage: it allows for the interpretation of the touching of the ABS spectrum with the continuum at $E=\Delta$ as a topological feature. Namely, for energy-independent scattering, we find the symmetry
@@
Evidently, this symmetry guarantees that for $\widehat{\phi}=0$ all eigenvalues of $h_{\text{eff}}$ go to 00, and thus $E=\pm\Delta$. The above statement is true for the here considered conventional s-wave superconducting contacts. Note that if we were to replace the contacts by topological p-wave superconductors, the Andreev reflection coefficient $\alpha$ would acquire an additional $\pi/2$ phase shift (see, e.g., Refs. [23, 26, 27]). This shift gives rise to the well-known correspondence of the ABS spectra between conventional and topological superconducting junctions, where energies $E$ for conventional superconductors map onto energies $\sqrt{\Delta^{2}-E^{2}}$ for topological superconductors (and vice versa). Consequently, for topological superconductors, the eigenvalues of $h_{\text{eff}}$ would return the actual ABS energies $E$ instead of $\sqrt{\Delta^{2}-E^{2}}$. By the same token, for topological superconductors, the very same chiral symmetry guarantees $E=0$ (due to the presence of Majorana zero energy states) instead of $E=\Delta$.
@@
Crucially, this correspondence ceases to be valid for the nonlinear case, that is, when $S_{e}$ is a function of $E$. This can be seen already in first order, $S_{e}(E)\approx S_{e}+E\partial S_{e}$, where the effective Hamiltonian (at $\widehat{\phi}=0$) simplifies to
$h_{\text{eff}}\approx-i\Delta\left(\begin{array}[]{cc}0&-E\partial S_{e}^{*}\\ E\partial S_{e}&0\end{array}\right)\ .$ (28)
@@
For the here considered conventional superconductors, a possible solution at $\widetilde{E}=0$ gets mapped to $E=\pm\Delta$, which has to be self-consistently reinserted into $h_{\text{eff}}$. But this nonlinear correction breaks the chiral symmetry, leading to a gapping of $\widetilde{E}$, or equivalently, to a detaching of $E$ from the continuous bands at ±$\pm\Delta$. Note that for topological superconductors, an energy-dependent scattering matrix does not lead to a similar breaking of the chiral symmetry, because $E$-dependent corrections to $S_{e}$ obviously vanish at $E=0$. For illustration purposes, we explicitly show this detaching mechanism in Figs. 2a (for $E$) and 2b (for $\widetilde{E}$). For concreteness, the energy spectra are computed with the explicit $L$-junction geometry, which is presented in Sec. VI.1. Details for the calculation of Fig. 2 can be found in Appendix A.
@@
Due to the above findings, one might think now that energy-dependent scattering inevitably causes a detaching to occur. However, it will turn out that in general this is not necessarily the case. By investigating various distinct origins of energy dependent scattering, we discover in what follows a mechanism by which the detaching is compensated for in certain limits. In particular, while a strong nonlinearity in the dispersion relation leads to a strongly energy-dependent scattering matrix, it, importantly, cannot open a gap at ${E=\Delta}$ (c.f. blue dotted curve in Figure 2). This is essentially due to the fact that the strong energy-dependence here does not come from a long dwelling time of the electrons inside the scattering region (i.e., the Thouless energy still remains large, $E_{\text{Th}}\rightarrow\infty$). While this finding may seem intuitively plausible, showing it rigorously on a mathematical level turns out to be a nontrivial task. In particular, in order to see the gap closing, and an effective restoring of chiral symmetry, one needs to go beyond the commonly made approximations (detailed in Sec. II). This is especially important for multi-channel conductors with strong cross-channel scattering.
@@
Figure 2: Impact of finite energy dependence in the scattering $S$ and nonzero backscattering in $R$, due to finite dispersion, on the ABS spectrum in a single-channel device. The spectrum shown above is obtained using the specific example of an $L$-junction geometry as discussed in detail in Sec. VI (see also Appendix A). Panels a) and b) show energy $E$ and pseudo-energy $\widetilde{E}$, respectively, for the same system parameters. With the standard framework employed – where energy dependence in $S$ is omitted and $R$ is applied as in Eq. (11) – the spectrum (black dotted curve) touches the gap ${E=\Delta}$ or ${\widetilde{E}=0}$, respectively, at ${\phi=0}$. When including energy dependent scattering $S\left[E\right]$ (while keeping the standard $R$) we observe a detaching from ${E=\Delta}$ and a gapping at ${\tilde{E}=0}$ in the pseudoenergy, respectively [green dashed curves in a) and b)]. If the energy-dependence in $S$ were due to a finite dwelling time, the green curve would be exact. Yet, for the parameters chosen here, the finite energy dependence is attributed to finite dispersive effects (whereas the dwelling time remains negligible). In this case, nonlinear dispersion effects in both the $R$- and $S$-matrices lead to a restoration of the gap closing (blue dotted curve), as argued in Sec. III.3.
@@ III.2 Finite dwelling time versus finite dispersion
$\psi_{x>L/2}$ $=T\left(E\right)e^{ik\left(E\right)x}\ .$ (30)
Since the central scattering region is arbitrary, we do not have a specific ansatz for the wave function for $-L/2<x<L/2$ (in fact, it is not required for the sake of our argument). The wave function, however, does include the possibility of a generic dispersion relation $k(E)$ and energy-dependent scattering coefficients, $R(E)$ and $T(E)$. Throughout this paper, we adopt the following language with regard to dispersion. If the dispersion relation for all relevant modes can be linearized, then we refer to this as the dispersion-free case, or to dispersion being absent (as wave packets here propagate over time without changing their width). If the linear approximation is invalid, we speak of a dispersive case, or finite dispersion.
To proceed, we expand all of these quantities up to first order, that is, $k(E)\approx k_{F}+E/v_{F}$  [^3] It might seem that expansion of $k(E)$ up to linear order is consistent with linear dispersion, but, as we will show shortly, this correction is enough to take into account the lowest order contributions to finite dispersion of wave packets. (in the single channel limit, we only have a single Fermi wave vector and Fermi velocity, $k_{F}$ and $v_{F}$) and $R(E)\approx R+E\partial R$ (likewise for $T$), where $E$ is still defined with respect to the chemical potential, such that this expansion corresponds to a Taylor series of $S$ in orders of $E$ around ${E=0}$.
We now initialize a Gaussian wave packet, centered around $E=0$,
[…]
Note that the factor $1/\sqrt{2\pi}\delta E$ is added only for convenience, and does not normalize the wave function (as normalization is not essential here). Up to linear order in $E$ this yields a wave function of the form
$+e^{-ik_{F}x}\left[R+iv_{F}\partial R\partial_{x}\right]g\left(x+v_{F}t\right),$ (32)
$\psi_{P}\left(x>L/2,t\right)$ $\approx e^{ik_{F}x}\left[T-iv_{F}\partial T\partial_{x}\right]g\left(x-v_{F}t\right),$ (33)
with the shortcut definition ${g\left(y\right)=\exp[-\delta E^{2}y^{2}/(2v_{F}^{2})]}$ for the Gaussian. We see that the linear corrections $\partial R$ and $\partial T$ give rise to a small deviation of the wave packet from a clean Gaussian (depending on the derivative of the Gaussian $\partial_{x}g$), which emerge upon impact of the wave packet at the scattering region (at around $t=0$). A revealing quantity to look at is the total overlap
[…]
Obviously, the sum $P_{S}+P_{\text{arms}}$ needs to stay constant due to probability conservation. Inserting Eqs. (32) and (33) into $P_{\text{arms}}$, and expanding in orders of $\partial R,\partial T$, $P_{\text{arms}}=P_{\text{arms}}^{(0)}+P_{\text{arms}}^{(1)}+\mathcal{O}[\partial R^{2},\partial T^{2},\partial R\partial T]$, we get in zeroth order
[…]
$P_{\text{arms}}^{(1)}=\frac{iv_{F}}{2}\left[\partial RR^{*}-R\partial R^{*}+\partial TT^{*}-T\partial T^{*}\right]g^{2}\left(v_{F}t\right)\ .$ (37)
We here present an important caveat to the above picture. Consider, for instance, a scattering region with infinitesimal size, such as a Dirac delta function potential as a well-documented textbook example, where the scattering coefficients depend explicitly on $k$, and thus, on $E$. Another example will be treated below: nanobridges with nontrivial device geometry. In all of these examples the electron spends manifestly negligible time inside the scattering region, $P_{S}\rightarrow 0$ (such that it is justified to assume the scattering size $L$ to be zero). Here, the above calculation is at risk of creating a conundrum in the form of a nonconserved probability, as the temporal “dip” in probability $\sim g^{2}(v_{F}t)$ in Eq. (37) would in this case go unaccounted for. But this looming contradiction can be resolved very simply by undoing an approximation we made when deriving Eq. (37): assuming $k_{F}$ infinitely large. Including instead a finite $k_{F}$, one can see that integrands of the form $\sim e^{\pm i2k_{F}x}g(x-v_{F}t)g(x+v_{F}t)$, which were previously neglected, lead to an important correction of the same order. Namely, we now find (assuming $\delta E\ll v_{F}k_{F}$)
$\begin{split}P_{\text{arms}}^{(1)}=\frac{iv_{F}}{2}\left[\frac{R-R^{*}}{v_{F}k_{F}}+\partial RR^{*}-R\partial R^{*}\right.\\ \left.+\partial TT^{*}-T\partial T^{*}\right]g^{2}\left(v_{F}t\right)\ .\end{split}$ (38)
$\frac{R-R^{*}}{v_{F}k_{F}}+\partial RR^{*}-R\partial R^{*}+\partial TT^{*}-T\partial T^{*}=0\ ,$ (39)
which thus relates the zeroth order scattering matrix elements, $R,T$, to its first order corrections, $\partial R,\partial T$, in a way which could not possibly be derived from unitarity of the scattering matrix alone (simply because it is less general, requiring $P_{\text{S}}=0$ as an additional assumption). Note that a similar condition can be derived, when including the state with an incoming wave from the right hand side,
[…]
$\frac{T-T^{*}}{v_{F}k_{F}}+\partial TR^{*}-T\partial R^{*}+\partial RT^{*}-R\partial T^{*}=0\ .$ (42)
$\frac{S_{e}-S_{e}^{*}}{2\mu}+\partial S_{e}S_{e}^{*}-S_{e}\partial S_{e}^{*}=0\ ,$ (43)
where we have used the identity $v_{F}k_{F}=2\mu$.
@@
Importantly, while for both a finite dwelling time (nonzero $P_{S}$ close to $t\approx 0$) or a finite $k_{F}$ (with $P_{S}=0$ for all $t$ due to $L=0$) the deviation from the Gaussian wave packet is the same ($\sim\partial_{x}g$), the physical origin could not be more different. In the first case, the probability of the electron to reside inside the conductor arms temporarily diminishes, whereas in the latter case it does not – leading to the extra requirement given in Eq. (43). Crucially, we will show further below, that if the latter is true, such that Eq. (43) applies, then the energy-dependence of the scattering matrix cannot open a gap at ${E=\Delta}$, and the chiral symmetry is preserved even for nonzero $\partial S_{e}$.
@@
As an intermediate comment, the reader might wonder, why this discussion is important, as it is common for solid state systems to assume $k_{F}$ large (thermodynamic limit). However, it is well-known that finite chemical potential effects are important for systems with low charge carrier density. In addition, our argument also matters for a system with a large electron density (in the solid state sense) in the multi-channel case as introduced in the dispersion relation given in Eq. (4). Notice that each channel has effectively a different chemical potential, due to the energy contribution of the transversal standing wave $\epsilon_{n}$. Hence, for a generic multi-channel situation, individual channels may have a noticeable dispersion especially when tuned to a regime where the number of occupied channels changes, as we will illustrate in more detail below.
@@ III.3 Closing the gap at $E=\Delta$
Up to now, we demonstrated that a finite energy-dependence of the scattering matrix does not necessarily indicate a finite dwelling time of the electrons inside the scattering region, but may also stem from a finite non-linearity in the dispersion relation of the leads. We now show how this fact enters in the behavior of the ABS spectrum close to the gap.
@@
The ordinary Andreev electron-hole receives a small correction ${\sim\frac{E}{2\mu}}$ , and there is in addition a finite electron-electron (respectively, hole-hole) backscattering, ${\sim\frac{\Delta}{2\mu}}$. We stress that the finite normal backscattering does not originate from an impurity at the $SN$-interface, but instead from a slight, but finite mismatch of the electron and hole wave vectors – a consequence of $\mu$ being finite  [^4] In alignment with Sec. II, unitarity only holds for $R$-matrix coefficients when normalizing with respect to the (mismatched) electron and hole group velocities.
@@
$h_{\text{eff}}=-i\Delta\left(\begin{array}[]{cc}\frac{\Delta}{4\mu}\left(S_{e}-S_{e}^{*}\right)&-E\partial S_{e}^{*}\\ E\partial S_{e}&\frac{\Delta}{4\mu}\left(S_{e}-S_{e}^{*}\right)\end{array}\right)\ ,$ (45)
@@
In what follows, we make two additional general statements. First, we argue that in the presence of multi-channel scattering, a strong energy-dependence may occur in the scattering matrix – yet again, not due to a finite dwelling time, but due to a variant of the same finite dispersive effects involving evanescent modes. Then, we argue why it is nonetheless difficult to generalize the above simple proof of the survival of the gap closing at $E=\Delta$ for $\phi=0$ by means of probability conservation arguments. Nonetheless, after updating the Beenakker formalism to explicitly take evanescent modes into account, we will be able to formulate a new type of constraint that, crucially, does not rely on probability conservation.

### v2

@@ III.1 Detaching from continuum
To set the stage, let us first consider a generic case of an energy-dependant scattering matrix, and show how this energy-dependence is noticeable within the ABS spectrum, at least within the above described standard literature formalism.
For this purpose, we focus in particular on the previously discussed fact that the entire ABS energy spectrum approaches the gap $\Delta$ for zero phase differences, ${\widehat{\phi}=0}$, and thus touches the continuum. As already explained, for a symmetric scattering matrix, the origin of the ${E=\Delta}$ solution can be conveniently seen on the level of Eq. (18), where for ${E=\Delta}$ and ${\widehat{\phi}=0}$, ${\alpha^{2}S_{e}e^{i\widehat{\phi}}S_{h}e^{-i\widehat{\phi}}=S_{e}S_{e}^{*}=1}$. If we now include a finite energy dependence in the scattering, we instead get the term $S_{e}(E)S_{e}^{*}(-E)$ (due to particle-hole symmetry), which in general no longer reduces to one. As a consequence, the ABS spectrum detaches from the continuum at $\widehat{\phi}=0$.
@@
The second peculiarity is that it does not return the eigenenergies directly, but instead, it provides what we refer to as the pseudoenergy $\widetilde{E}=\sqrt{\Delta^{2}-E^{2}}$. This is because the subtraction of Eq. (23) from Eq. (22) yields $\text{Im}[\alpha]=-\sqrt{1-E^{2}/\Delta^{2}}$ on the right hand side. Obviously, the addition of Eqs. (22) and (23) (which was the original procedure of Ref. [13]) would instead yield $\text{Re}[\alpha]=E/\Delta$ and would thus be much closer to the generally more conventional form of an effective Schrödinger equation. But, the form chosen above has a central advantage: it allows for the interpretation of the touching of the ABS spectrum with the continuum at $E=\Delta$ as a topological feature. Namely, for energy-independent scattering, we find the symmetry
@@
Evidently, this symmetry guarantees that for $\widehat{\phi}=0$ all eigenvalues of $h_{\text{eff}}$ go to zero, and thus $E=\pm\Delta$. The above statement is true for the here considered conventional s-wave superconducting contacts. Note that if we were to replace the contacts by topological p-wave superconductors, the Andreev reflection coefficient $\alpha$ would acquire an additional $\pi/2$ phase shift (see, e.g., Refs. [23, 26, 27]). This shift gives rise to the well-known correspondence of the ABS spectra between conventional and topological superconducting junctions, where energies $E$ for conventional superconductors map onto energies $\sqrt{\Delta^{2}-E^{2}}$ for topological superconductors (and vice versa). Consequently, for topological superconductors, the eigenvalues of $h_{\text{eff}}$ would return the actual ABS energies $E$ instead of $\sqrt{\Delta^{2}-E^{2}}$. By the same token, for topological superconductors, the very same chiral symmetry guarantees $E=0$ (due to the presence of Majorana zero energy states) instead of $E=\Delta$.
@@
Crucially, this correspondence ceases to be valid for the nonlinear case, that is, when $S_{e}$ is a function of $E$. This can be seen already in first order, $S_{e}(E)\approx S_{e}+E\delta S_{e}$, where the effective Hamiltonian (at $\widehat{\phi}=0$) simplifies to
$h_{\text{eff}}\approx-i\Delta\left(\begin{array}[]{cc}0&-E\delta S_{e}^{*}\\ E\delta S_{e}&0\end{array}\right)\ .$ (28)
@@
For the here considered conventional superconductors, a possible solution at ${\widetilde{E}=0}$ gets mapped to ${E=\pm\Delta}$, which has to be self-consistently reinserted into $h_{\text{eff}}$. But this nonlinear correction breaks the chiral symmetry, leading to a gapping of $\widetilde{E}$, or equivalently, to a detaching of $E$ from the continuous bands at ±$\pm\Delta$. Note that for topological superconductors, an energy-dependant scattering matrix does not lead to a similar breaking of the chiral symmetry, because $E$-dependant corrections to $S_{e}$ obviously vanish at ${E=0}$. For illustration purposes, we explicitly show this detaching mechanism in Figs. 2a (for $E$) and 2b (for $\widetilde{E}$). For concreteness, the energy spectra are computed with the explicit $L$-junction geometry, which is presented in Sec. C. Details for the calculation of Fig. 2 can be found in Appendix A.
@@
Due to the above findings, one might think now that energy-dependant scattering inevitably causes a detaching to occur. However, it will turn out that in general this is not necessarily the case. By investigating various distinct origins of energy dependant scattering, we discover in what follows a mechanism by which the detaching is compensated for in certain limits. In particular, while a strong nonlinearity in the dispersion relation leads to a strongly energy-dependant scattering matrix, it, importantly, cannot open a gap at ${E=\Delta}$ (cf. blue dotted curve in Figure 2). This is essentially due to the fact that the strong energy-dependence here does not come from a long dwelling time of the electrons inside the scattering region (i.e., the Thouless energy still remains large, $E_{\text{Th}}\rightarrow\infty$). While this finding may seem intuitively plausible, showing it rigorously on a mathematical level turns out to be a nontrivial task. In particular, in order to see the gap closing, and an effective restoring of chiral symmetry, one needs to go beyond the commonly made approximations (detailed in Sec. II). This is especially important for multi-channel conductors with strong cross-channel scattering.
@@
Figure 2: Impact of finite energy dependence in the scattering $S$ and nonzero backscattering in $R$, due to finite dispersion, on the ABS spectrum in a single-channel device. The spectrum shown above is obtained using the specific example of an $L$-junction geometry as discussed in detail in Sec. VI (see also Appendix A). Panels a) and b) show the energy $E$ and pseudo-energy $\widetilde{E}$, respectively, for the same system parameters. With the standard framework employed – where energy dependence in $S$ is omitted and $R$ is applied as in Eq. (11) – the spectrum (black dotted curve) touches the gap ${E=\Delta}$ or ${\widetilde{E}=0}$, respectively, at ${\phi=0}$. When including energy dependant scattering $S\left[E\right]$ (while keeping the standard $R$) we observe a detaching from ${E=\Delta}$ and a gapping at ${\tilde{E}=0}$ in the pseudoenergy, respectively [green dashed curves in a) and b)]. If the energy-dependence in $S$ were due to a finite dwelling time, the green curve would be exact. Yet, for the parameters chosen here, the finite energy dependence is attributed to finite dispersive effects (whereas the dwelling time remains negligible). In this case, nonlinear dispersion effects in both the $R$- and $S$-matrices lead to a restoration of the gap closing (blue dotted curve), as argued in Sec. III.3.
@@ III.2 Finite dwelling time versus finite dispersion
Figure 3: Probability conservation at the time of scattering, depending on the size of the scattering region or on the presence of evanescent modes. A Gaussian wave packet is prepared and impinging upon a scattering centre. Unitarity of the $S$-matrix ensures probability conservation before ($t_{1}$) and after ($t_{3}$) the scattering in the conductor arms. However, during the scattering process (around $t_{2}$), properties beyond unitarity become visible, depending on the situation. If the scattering centre has non-negligible width [panel a)], part of the probability current flows into the scattering region and is temporarily lost for the conductor arms (leading to a finite energy-dependence of $S$). If we consider instead the case of cross-channel scattering with dispersive evanescent modes [blue real exponentials in panel b)], probability current temporarily flows into these virtually occupied higher channels, even if the scattering region is short (negligible dwelling time ${\tau_{\mathrm{dw}}=1/E_{\mathrm{Th}}\to 0}$). In this latter case, the finite energy-dependence of the planar part of $S$ is compensated through a constraint related to planar-evanescent (and evanescent-planar) scattering processes, cf. Eq. (66) (one of the central results of this work).
Typically, probability conservation is expressed by the unitarity of the scattering matrix which aggregates and relates the total probability densities well before and after the time of scattering ($t_{1}$ and $t_{3}$ in Fig. 3). More generally, however, probability has to be conserved at all times, particularly during the scattering process itself (around $t_{2}$). By tracking the time-evolution of Gaussian wave packets and integrating over the probability densities, we find that under certain circumstances, there emerge conditions beyond unitarity. The derivation of these conditions is detailed in Appendix B. Below we summarise the results. We consider a single channel scattering problem in two particular limits.
In a first case, we assume that the scattering centre has a finite size $L$ which, during the actual scattering process, temporarily absorbs a portion of the probability density. This results in a distortion of the Gaussians, the magnitude of which can be related to a term ${\sim\delta S_{\mathrm{e}}S_{\mathrm{e}}^{*}-S_{\mathrm{e}}\delta S_{\mathrm{e}}^{*}}$, where $\delta S_{\mathrm{e}}$ captures the linear energy dependence of $S_{\mathrm{e}}$. This case corresponds to the view usually taken in the existing literature, where a finite dwelling time gives rise to a finite energy dependence of the scattering matrix.
We contrast the above with a second case with two important differences. On the one hand, we assume that the scattering region has negligible width ${L=0}$, such that the above argument no longer applies. Conversely, however, we now assume that the momentum $k\left(E\right)$ can no longer be linearized. Throughout this paper, we refer to this as the “dispersive” case (or to a regime of “finite dispersion”). We now still have a finite energy dependence in the scattering problem, this time stemming from a finite dispersion instead of a finite size of the scattering region. This results first of all in an additional non-Gaussian correction term ${\sim S_{\mathrm{e}}-S_{\mathrm{e}}^{*}}$. On the other hand, due to the negligible scattering size, we still have to impose probability conservation everywhere within the conductor arms, leading to an additional condition with respect to the electron scattering matrix $S_{\mathrm{e}}$ as
$\frac{S_{\mathrm{e}}-S_{\mathrm{e}}^{*}}{2\mu}+\delta S_{\mathrm{e}}S_{\mathrm{e}}^{*}-S_{\mathrm{e}}\delta S_{\mathrm{e}}^{*}=0\ .$ (29)
@@
We emphasise that while for both cases, the deviation from the Gaussian wave packet is of the exact same shape (cf. Appendix B), the physical origin could not be more different. In the first case, the probability of the electron to reside inside the conductor arms temporarily diminishes, whereas in the latter case it does not – leading to the extra requirement given in Eq. (29). Crucially, we show further below, that if the latter is true, such that Eq. (29) applies, the energy-dependence of the scattering matrix cannot open a gap at ${E=\Delta}$, and the chiral symmetry is preserved even for nonzero $\delta S_{e}$.
@@
As an intermediate comment, the reader might wonder, why this discussion is important, as it is common for solid state systems to assume $k_{F}$ large (thermodynamic limit). However, it is well-known that finite chemical potential effects are important for systems with low charge carrier density. In addition, our argument also matters for a system with a large electron density (in the solid state sense) in the multi-channel case as introduced in the dispersion relation given in Eq. (4). Notice that each channel has effectively a different chemical potential, due to the energy contribution of the transversal standing wave $\epsilon_{n}$. Hence, for a generic multi-channel situation, individual channels may have a noticeable dispersion especially when tuned to a regime where the number of occupied channels changes, as we illustrate in more detail below.
@@ III.3 Closing the gap at $E=\Delta$
Up to now, we demonstrated that a finite energy-dependence of the scattering matrix does not necessarily indicate a finite dwelling time of the electrons inside the scattering region, but may also stem from a finite non-linearity in the dispersion relation of the leads. We now show how this fact enters in the behaviour of the ABS spectrum close to the gap.
@@
The ordinary Andreev electron-hole receives a small correction ${\sim\frac{E}{2\mu}}$ , and there is in addition a finite electron-electron (respectively, hole-hole) backscattering, ${\sim\frac{\Delta}{2\mu}}$. We stress that the finite normal backscattering does not originate from an impurity at the $SN$-interface, but instead from a slight, but finite mismatch of the electron and hole wave vectors – a consequence of $\mu$ being finite [^3] In alignment with Sec. II, unitarity only holds for $R$-matrix coefficients when normalising with respect to the (mismatched) electron and hole group velocities..
@@
$h_{\text{eff}}=-i\Delta\left(\begin{array}[]{cc}\frac{\Delta}{4\mu}\left(S_{e}-S_{e}^{*}\right)&-E\delta S_{e}^{*}\\ E\delta S_{e}&\frac{\Delta}{4\mu}\left(S_{e}-S_{e}^{*}\right)\end{array}\right)\ ,$ (31)
@@
In what follows, we make two additional general statements. First, we argue that in the presence of multi-channel scattering, a strong energy-dependence may occur in the scattering matrix – yet again, not due to a finite dwelling time, but due to a variant of the same finite dispersive effects involving evanescent modes. Then, we argue why it is nonetheless difficult to generalise the above simple proof of the survival of the gap closing at $E=\Delta$ for $\phi=0$ by means of probability conservation arguments. Nonetheless, after updating the Beenakker formalism to explicitly take evanescent modes into account, we can formulate a new type of constraint that, crucially, does not rely on probability conservation.